OpenStudy (anonymous):

-9x + 7 = 25 + 2(5 - x)

OpenStudy (anonymous):

x= -4 -9x + 7 = 25 = 2(5 - x) -9x +7 = 25 + 10 - 2x -9x + 28 - 2x -7x = 28 x = 28/-7 x = -4